OpenStudy (christos):

Given that sinƟ = sqrt(7)/4 and cosƟ = 3/4 find the exact value of tag = Ɵ

OpenStudy (anonymous):

\[\tan(\theta)=\frac{\sin(\theta)}{\cos(\theta)}\]

OpenStudy (anonymous):

\[\frac{\frac{\sqrt{7}}{4}}{\frac{3}{4}}\] the fours cancel

OpenStudy (christos):

hmmm ill try and let you on